### v1

@@ Abstract
A singularly perturbed reaction-diffusion problem posed on the unit square in $\mathbb{R}^{2}$ is solved numerically by a local discontinuous Galerkin (LDG) finite element method. Typical solutions of this class of problem exhibit boundary layers along the sides of the domain; these layers generally cause difficulties for numerical methods. Our LDG method handles the boundary layers by using a Shishkin mesh and also introducing the new concept of a “layer-upwind flux”—a discrete flux whose values are chosen on the fine mesh (which lies inside the boundary layers) in the direction where the layer weakens. On the coarse mesh, one can use a standard central flux. No penalty terms are needed with these fluxes, unlike many other variants of the LDG method. Our choice of discrete flux makes it feasible to derive an optimal-order error analysis in a balanced norm; this norm is stronger than the usual energy norm and is a more appropriate measure for errors in computed solutions for singularly perturbed reaction-diffusion problems. It will be proved that the LDG method is usually convergent of order $O((N^{-1}\ln N)^{k+1})$ in the balanced norm, where $N$ is the number of mesh intervals in each coordinate direction and tensor-product piecewise polynomials of degree $k$ in each coordinate variable are used in the LDG method. This result is the first of its kind for the LDG method applied to this class of problem and is optimal for convergence on a Shishkin mesh. Its sharpness is confirmed by numerical experiments.
@@ 1 Introduction
$u$ $=0\ \text{ on }\partial\Omega,$ (1.1b)
@@ 1.2 The LDG method
The local discontinuous Galerkin (LDG) method [13] is a popular FEM that uses discontinuous piecewise polynomial spaces for the trial and test solutions. Its basic idea is to rewrite the PDE as an equivalent first-order system to which one then applies the original DG method [29]. In the LDG method one must specify a numerical flux on the boundary of each mesh element; if this is done appropriately, then the auxiliary variables that approximate the derivatives of the solution can be eliminated locally, which explains the name LDG. The LDG method is superior to many other numerical methods because of its strong stability, high-order accuracy, flexibility for $hp$-adaptivity, and high parallelizability. It is well suited to problems whose solutions have boundary layers. Furthermore, it can produce high-order accuracy for both the solution itself and its gradient; the latter quantity is important in fluid dynamics applications where values of pressure gradients are desired.
@@ 1.4 Complexity of the error analysis
It is notable that our new error analysis on the fine Shishkin mesh makes no use of the nodal superconvergence property of the $L^{2}$ projection on locally uniform meshes that was mentioned in Section 1.3, thereby opening the possibility of extending this work to other layer-adapted meshes such as Bakhvalov-type meshes. In general, our new layer-upwinded numerical flux appears to be a very promising tool in the application of the LDG method to other problems whose solutions have boundary layers.
@@ 2.1 Problem and solution properties
If for instance one assumes that $f,b\in C^{4,\alpha}(\bar{\Omega})$ and the corner compatibility conditions $f(0,0)=f(1,0)=f(0,1)=f(1,1)=0$ are satisfied, then Lemma  2.1 holds true for $m=2$; see [25, Section 1.2].
@@ 2.2 Shishkin mesh
Let $N$ be a positive integer that is divisible by $4$. Our mesh uses $N+1$ points in each coordinate direction. Define the mesh transition parameter $\tau$ by
@@
Figure 1 displays a Shishkin mesh with $N=8$. It is uniform and coarse on the region $\Omega_{22}:=(\tau,1-\tau)\times(\tau,1-\tau)$, but is refined in the layer regions
@@ 2.3 The LDG method with layer-upwind flux
The LDG finite element method has been used to solve many types of problem, including singularly perturbed elliptic problems; see [7] and its bibliography. Our construction of it here is standard—apart from the choice of numerical fluxes in (2.9), which is new and is crucial for our error analysis.
@@
Note that our LDG method does not use any penalty terms when defining the flux for singularly perturbed problems, unlike many other LDG papers [7, 9, 10, 11, 40].
A compact form of the LDG method is got by summing (2.8) over all $K_{ij}\in\Omega_{N}$: Find $\bm{\mathbbm{W}}=(\mathbbm{U},\mathbbm{P},\mathbbm{Q})\in\mathcal{V}_{N}^{3}$ such that $B(\bm{\mathbbm{W}};\bm{\mathbbm{z}})=\left\langle f,\mathbbm{v}\right\rangle$ for all $\bm{\mathbbm{z}}=(\mathbbm{v},\mathbbm{s},\mathbbm{r})\in\mathcal{V}_{N}^{3}$, where
@@
$\interleave\bm{\mathbbm{z}}\interleave_{B}^{2}$ $=\varepsilon^{-3/2}\left(\left\|\mathbbm{s}\right\|^{2}+\left\|\mathbbm{r}\right\|^{2}\right)+\left\|b^{1/2}\mathbbm{v}\right\|^{2}$ (2.13a)
@@ 3.2 Stability and approximation properties of projectors
Let
$\mathcal{R}\in\{\Pi,\Pi_{x}^{-},\Pi_{x}^{+},\Pi_{y}^{-},\Pi_{y}^{+},\Pi_{xy}^{+},\Pi_{xy}^{\pm},\Pi_{xy}^{\mp},\Pi_{xy}^{-}\}$. Similarly to [8, Lemma 5], one has the following stability properties in the $L^{\infty}$ and $L^{2}$ norms (here $C>0$ is some generic constant as described at the end of Section 1):
@@
$\left\|\Pi^{-}_{xy}z\right\|_{K_{ij}}$ $\leq C\Big{[}\left\|z\right\|_{K_{ij}}+h^{1/2}_{j}\left\|z^{-}_{x,j}\right\|_{I_{i}}+h^{1/2}_{i}\left\|z^{-}_{i,y}\right\|_{J_{j}}+h^{1/2}_{i}h^{1/2}_{j}|z(x_{i}^{-},y_{j}^{-})|\Big{]}.$ (3.4e)
Analogous bounds are valid for the remaining projectors $\Pi_{x}^{+},\Pi_{y}^{+},\Pi_{xy}^{\pm},\Pi_{xy}^{\mp}$ and $\Pi_{xy}^{+}$. Furthermore, one has the following approximation properties (see, e.g., [8, Lemma 3] and [40, Lemma 4.3] for the case $k>0$, and [17, Lemma 3.8] and [18, p.60] for the case $k=0$):
@@
Assume that Lemma 2.1 is valid for $m=k$. Then for the solution $u$ of (1.1), there exists a constant $C>0$ such that
@@
$\leq C\left\{h^{k+1}\left[\int_{K_{ij}}\varepsilon^{-(k+1)}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\,\mathrm{d}y\right]^{1/2}+H^{k+1}\left[\int_{K_{ij}}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\,\mathrm{d}y\right]^{1/2}\right\}$
@@
$\leq C\left\{\left[\int_{K_{ij}}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\,\mathrm{d}y\right]^{1/2}+h_{j}^{1/2}\left[\int_{I_{i}}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\right]^{1/2}\right\}$
@@
$\left\|\eta_{u^{\mathrm{b}}_{1}}\right\|_{\Omega_{21}}$ $\leq C\left(\int_{0}^{\tau}\int_{\tau}^{1-\tau}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\,\mathrm{d}y\right)^{1/2}\leq C\tau^{1/2}\varepsilon^{1/4}N^{-\sigma}\leq C\varepsilon^{1/2}N^{-\sigma}(\ln N)^{1/2}.$ (3.9)
@@
$\left\|\eta_{u^{\mathrm{b}}_{1}}\right\|_{\Omega_{23}}\leq C\varepsilon^{1/4}N^{-\sigma}(\ln N)^{1/2}.$ (3.10)
@@
$\left\|\eta_{u^{\mathrm{b}}_{1}}\right\|_{\Omega_{22}}\leq\left\|u^{\mathrm{b}}_{1}\right\|_{\Omega_{22}}\leq\left\|e^{-\beta x/\sqrt{\varepsilon}}\right\|_{\Omega_{22}}\leq C\varepsilon^{1/4}N^{-\sigma}.$ (3.11)
@@
If $K_{ij}\in\Omega_{12}$, then $\eta_{u^{\mathrm{c}}_{1}}=u^{\mathrm{c}}_{1}-\Pi_{x}^{-}u^{\mathrm{c}}_{1}$; use the $L^{2}$-stability property (3.4d) and the bound (2.1c) to obtain
@@
$\leq C\left\{\left[\int_{K_{ij}}e^{-2\beta(x+y)/\sqrt{\varepsilon}}\,\mathrm{d}x\,\mathrm{d}y\right]^{1/2}+h_{i}^{1/2}\left[\int_{J_{j}}e^{-2\beta x_{i}/\sqrt{\varepsilon}}e^{-2\beta y/\sqrt{\varepsilon}}\,\mathrm{d}y\right]^{1/2}\right\}$
$\leq Ch^{1/2}\left\|e^{-\beta y/\sqrt{\varepsilon}}\right\|_{J_{j}},$
@@
$\left\|\eta_{u^{\mathrm{c}}_{1}}\right\|_{\Omega_{22}}\leq\left\|u^{\mathrm{c}}_{1}\right\|_{\Omega_{22}}\leq\left\|e^{-\beta(x+y)/\sqrt{\varepsilon}}\right\|_{\Omega_{22}}\leq C\varepsilon^{1/2}N^{-2\sigma}.$ (3.15)
@@
$\left\|\eta_{u^{\mathrm{c}}_{1}}\right\|_{\Omega\backslash\Omega_{22}}\leq C\varepsilon^{1/4}(N^{-1}\ln N)^{k+1}\quad\text{and}\quad\left\|\eta_{u^{\mathrm{c}}_{1}}\right\|_{\Omega_{22}}\leq C\varepsilon^{1/2}N^{-2\sigma}.$ (3.16)
@@
$\left\|\eta_{u^{\mathrm{b}}_{1}}\right\|_{L^{\infty}(K_{ij})}\leq\left\|u^{\mathrm{b}}_{1}\right\|_{L^{\infty}(K_{ij})}\leq\left\|e^{-\beta x/\sqrt{\varepsilon}}\right\|_{L^{\infty}(K_{ij})}\leq CN^{-\sigma}.$ (3.19)
@@
$\left\|\eta_{u^{\mathrm{c}}_{1}}\right\|_{L^{\infty}(K_{ij})}\leq\left\|u^{\mathrm{c}}_{1}\right\|_{L^{\infty}(K_{ij})}\leq\left\|e^{-\beta(x+y)/\sqrt{\varepsilon}}\right\|_{L^{\infty}(K_{ij})}\leq CN^{-\sigma}.$ (3.21)
@@ 4 Superapproximation properties of projectors
In Section 4 we derive various superapproximation properties for the three types of vertex-edge-element approximations appearing in (3.3).
@@ 4.1 Superapproximation properties of element projector $\Pi$
We prove only the first inequality as the second is derived similarly. For $m=0,1,\dots$, let $\ell_{m}(x)$ be the rescaled Legendre polynomial of degree $m$ defined via Rodrigues’ formula for the standard Legendre polynomial ^$\hat{\ell}_{m}(s)$ on the interval $-1\leq s\leq 1$:
@@
By (4.3), (4.1), a scaling argument and integration by parts, one has
@@
$=\left|\frac{1}{2^{k+1+n}(k+1)!n!}\int_{\hat{K}}\hat{z}(s,t)\partial_{s}^{k+1}(s^{2}-1)^{k+1}\partial_{t}^{n}(t^{2}-1)^{n}\,\mathrm{d}s\,\mathrm{d}t\right|$
$=\left|\frac{1}{2^{k+1+n}(k+1)!n!}\int_{\hat{K}}[\partial_{s}^{k+1}\hat{z}(s,t)](s^{2}-1)^{k+1}\partial_{t}^{n}(t^{2}-1)^{n}\,\mathrm{d}s\,\mathrm{d}t\right|$ (4.6)
$\leq\frac{C}{2^{k+1}(k+1)!}\int_{\hat{K}}\left|\partial_{s}^{k+1}\hat{z}(s,t)\right|(s^{2}-1)^{k+1}\,\mathrm{d}s\,\mathrm{d}t$
$\leq C\left[\int_{\hat{K}}\left(\partial_{s}^{k+1}\hat{z}\right)^{2}\,\mathrm{d}s\,\mathrm{d}t\right]^{1/2}\leq C\left[\frac{1}{|K_{ij}|}\int_{K_{ij}}\left(h_{i}^{k+1}\partial_{x}^{k+1}z\right)^{2}\,\mathrm{d}x\,\mathrm{d}y\right]^{1/2}$
$\leq C\sqrt{\frac{h_{i}}{h_{j}}}\,h_{i}^{k}\left\|\partial_{x}^{k+1}z\right\|_{K_{ij}}.$
Similarly,
$|\alpha^{i+1,j}_{k+1,n}|\leq C\sqrt{\frac{h_{i+1}}{h_{j}}}\,h_{i+1}^{k}\left\|\partial_{x}^{k+1}z\right\|_{K_{i+1,j}}.$
Our argument now depends on the parity of the nonnegative integer $k$. If $k$ is odd, then $\alpha^{ij}_{k+1,n}+\alpha^{i+1,j}_{k+1,n}(-1)^{k+1}=\alpha^{ij}_{k+1,n}+\alpha^{i+1,j}_{k+1,n}$. Thus, since $h_{i}=h_{i+1}$, the pairwise orthogonality of the rescaled Legendre polynomials and Cauchy-Schwarz and inverse inequalities yield
@@
$\leq Ch_{j}^{1/2}\sqrt{\frac{h_{i}}{h_{j}}}\,h_{i}^{k}\left\|\partial_{x}^{k+1}z\right\|_{K_{ij}\cup K_{i+1,j}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$

### v3

@@ Abstract
A singularly perturbed reaction-diffusion problem posed on the unit square in $\mathbb{R}^{2}$ is solved numerically by a local discontinuous Galerkin (LDG) finite element method. Typical solutions of this class of 2D problems exhibit boundary layers along the sides of the domain; these layers generally cause difficulties for numerical methods. Our LDG method handles the boundary layers by using a Shishkin mesh and also introducing the new concept of a “layer-upwind flux”—a discrete flux whose values are chosen on the fine mesh (which lies inside the boundary layers) in the direction where the layer weakens. On the coarse mesh, one can use a standard central flux. No penalty terms are needed with these fluxes, unlike many other variants of the LDG method. Our choice of discrete flux makes it feasible to derive an optimal-order error analysis in a balanced norm; this norm is stronger than the usual energy norm and is a more appropriate measure for errors in computed solutions for singularly perturbed reaction-diffusion problems. It will be proved that the LDG method is usually convergent of order $O((N^{-1}\ln N)^{k+1})$ in the balanced norm, where $N$ is the number of mesh intervals in each coordinate direction and tensor-product piecewise polynomials of degree $k$ in each coordinate variable are used in the LDG method. This result is the first of its kind for the LDG method applied to this class of problem and is optimal for convergence on a Shishkin mesh. Its sharpness is confirmed by numerical experiments.
@@ 1 Introduction
$u$ $=0\hskip 28.45274pt\text{ on }\partial\Omega,$ (1.1b)
@@ 1.2 The LDG method
The local discontinuous Galerkin (LDG) method [13] is a popular FEM that uses discontinuous piecewise polynomial spaces for the trial and test solutions. Its basic idea is to rewrite the second-order differential equation as an equivalent first-order system to which one then applies the original DG method [29]. In the LDG method one must specify a numerical flux on the boundary of each mesh element; if this is done appropriately, then the auxiliary variables that approximate the derivatives of the solution can be eliminated locally, which explains the name LDG. The LDG method is superior to many other numerical methods because of its strong stability, high-order accuracy, flexibility for $hp$-adaptivity, and high parallelizability. It is well suited to problems whose solutions have boundary layers. Furthermore, it can produce high-order accuracy for both the solution itself and its gradient; the latter quantity is important in fluid dynamics applications where values of pressure gradients are desired.
@@ 1.4 Complexity of the error analysis
It is notable that our new error analysis on the fine Shishkin mesh makes no use of the nodal superconvergence property of the $L^{2}$ projection on locally uniform meshes that was mentioned in Section 1.3, thereby opening the possibility of extending this work to other layer-adapted meshes such as Shishkin-type and Bakhvalov-type meshes. In general, our new layer-upwinded numerical flux appears to be a very promising tool in the application of the LDG method to other problems whose solutions have boundary layers.
@@ 2.1 Problem and solution properties
If for instance one assumes that $f,b\in C^{4,\alpha}(\bar{\Omega})$ and the corner compatibility conditions $f(0,0)=f(1,0)=f(0,1)=f(1,1)=0$ are satisfied, then Lemma 2.1 holds true for $m=2$; see [25, Section 1.2].
@@ 2.2 Shishkin mesh
Let $N\geq 4$ be a positive integer that is divisible by $4$. Our mesh uses $N+1$ points in each coordinate direction. Define the mesh transition parameter $\tau$ by
@@
The right-hand picture of Figure 1 displays a Shishkin mesh with $N=8$. It is uniform and coarse on the region $\Omega_{22}:=(\tau,1-\tau)\times(\tau,1-\tau)$, but is refined in the layer regions
@@ 2.3 The LDG method with layer-upwind flux
Note that our LDG method has a certain degree of symmetry; furthermore, our method does not use any penalty terms when defining the flux for singularly perturbed problems, unlike many other LDG papers [7, 9, 10, 11, 40].
To write the LDG method in a compact form, sum (2.8) over all $K_{ij}\in\Omega_{N}$: Find $\bm{\mathbbm{W}}=(\mathbbm{U},\mathbbm{P},\mathbbm{Q})\in\mathcal{V}_{N}^{3}$ such that $B(\bm{\mathbbm{W}};\bm{\mathbbm{z}})=\left\langle f,\mathbbm{v}\right\rangle$ for all $\bm{\mathbbm{z}}=(\mathbbm{v},\mathbbm{s},\mathbbm{r})\in\mathcal{V}_{N}^{3}$, where
@@
$\interleave\bm{\mathbbm{z}}\interleave_{B}^{2}$ $=\varepsilon^{-3/2}\left(\left\|\mathbbm{s}\right\|^{2}+\left\|\mathbbm{r}\right\|^{2}\right)+\left\|b^{1/2}\mathbbm{v}\right\|^{2}.$ (2.13a)
@@ 3.2 Stability and approximation properties of projectors
Let $\mathcal{R}\in\{\Pi,\Pi_{x}^{-},\Pi_{x}^{+},\Pi_{y}^{-},\Pi_{y}^{+},\Pi_{xy}^{+},\Pi_{xy}^{\pm},\Pi_{xy}^{\mp},\Pi_{xy}^{-}\}$. Similarly to [8, Lemma 5], one has the following stability properties in the $L^{\infty}$ and $L^{2}$ norms (here $C>0$ is some generic constant as described at the end of Section 1):
@@
$\left\|\Pi^{-}_{xy}z\right\|_{K_{ij}}$ $\leq C\Big{[}\left\|z\right\|_{K_{ij}}+h^{1/2}_{j}\left\|z^{-}_{x,j}\right\|_{I_{i}}+h^{1/2}_{i}\left\|z^{-}_{i,y}\right\|_{J_{j}}+(h_{i}h_{j})^{1/2}|z^{-}_{ij}|\Big{]},$ (3.4e)
where $z^{-}_{ij}:=z(x_{i}^{-},y_{j}^{-})$. Analogous bounds are valid for the remaining projectors $\Pi_{x}^{+}$, $\Pi_{y}^{+}$, $\Pi_{xy}^{\pm}$, $\Pi_{xy}^{\mp}$ and $\Pi_{xy}^{+}$. Furthermore, one has the following approximation properties (see, e.g., [8, Lemma 3] and [40, Lemma 4.3] for the case $k>0$, and [17, Lemma 3.8] and [18, p.60] for the case $k=0$):
@@
Assume that Lemma 2.1 is valid for $m=k$. Then for the solution $(u,p,q)$ of (2.7), there exists a constant $C>0$ such that
@@
$\leq C\left\{h^{k+1}\Big{[}\int_{K_{ij}}\varepsilon^{-(k+1)}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\,\mathrm{d}y\Big{]}^{1/2}+H^{k+1}\Big{[}\int_{K_{ij}}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\,\mathrm{d}y\Big{]}^{1/2}\right\}$
@@
$\leq C\left\{\Big{[}\int_{K_{ij}}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\,\mathrm{d}y\Big{]}^{1/2}+h_{j}^{1/2}\Big{[}\int_{I_{i}}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\Big{]}^{1/2}\right\}$
@@
$\left\|\eta_{u^{\mathrm{b}}_{1}}\right\|_{\Omega_{21}}$ $\leq C\left\|e^{-\beta x/\sqrt{\varepsilon}}\right\|_{\Omega_{21}}\leq C\tau^{1/2}\left(\int_{\tau}^{1-\tau}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\right)^{1/2}$
$=C\tau^{1/2}\Big{[}\frac{\sqrt{\varepsilon}}{2\beta}e^{-2\beta\tau/\sqrt{\varepsilon}}(1-e^{-2\beta(1-2\tau)/\sqrt{\varepsilon}})\Big{]}^{1/2}$
$\leq C\tau^{1/2}\varepsilon^{1/4}N^{-\sigma}\leq C\varepsilon^{1/2}N^{-\sigma}(\ln N)^{1/2},$ (3.9)
where we used $\tau=\frac{\sigma\sqrt{\varepsilon}}{\beta}\ln N\leq 1/4$ implying $1-2\tau\geq 2\tau$ and $e^{-2\beta(1-2\tau)/\sqrt{\varepsilon}}\leq e^{-4\beta\tau/\sqrt{\varepsilon}}=N^{-4\sigma}\leq 4^{-4\sigma}\leq 1/16$ for $N\geq 4$ and $\sigma\geq\hat{k}+1\geq 1$.
@@
$\left\|\eta_{u^{\mathrm{b}}_{1}}\right\|_{\Omega_{23}}\leq C\varepsilon^{1/2}N^{-\sigma}(\ln N)^{1/2}.$ (3.10)
@@
$\left\|\eta_{u^{\mathrm{b}}_{1}}\right\|_{\Omega_{22}}\leq 2\left\|u^{\mathrm{b}}_{1}\right\|_{\Omega_{22}}\leq C\left\|e^{-\beta x/\sqrt{\varepsilon}}\right\|_{\Omega_{22}}\leq C\left(\int_{\tau}^{1-\tau}e^{-2\beta x/\sqrt{\varepsilon}}\,\mathrm{d}x\right)^{1/2}\leq C\varepsilon^{1/4}N^{-\sigma}.$ (3.11)
@@
If $K_{ij}\in\Omega_{12}$, then $\eta_{u^{\mathrm{c}}_{1}}=u^{\mathrm{c}}_{1}-\Pi_{x}^{-}u^{\mathrm{c}}_{1}$; use the $L^{2}$-stability property (3.4c) and the bound (2.1c) to obtain
@@
$\leq C\left[\left\|e^{-\beta(x+y)/\sqrt{\varepsilon}}\right\|_{K_{ij}}+h_{i}^{1/2}\left\|e^{-\beta(x_{i}+y)/\sqrt{\varepsilon}}\right\|_{J_{j}}\right]$
$\leq Ch_{i}^{1/2}\left\|e^{-\beta y/\sqrt{\varepsilon}}\right\|_{J_{j}},$
@@
$\left\|\eta_{u^{\mathrm{c}}_{1}}\right\|_{\Omega_{22}}\leq 2\left\|u^{\mathrm{c}}_{1}\right\|_{\Omega_{22}}\leq C\left\|e^{-\beta(x+y)/\sqrt{\varepsilon}}\right\|_{\Omega_{22}}\leq C\varepsilon^{1/2}N^{-2\sigma}.$ (3.15)
@@
$\left\|\eta_{u^{\mathrm{c}}_{1}}\right\|_{\Omega\backslash\Omega_{22}}\leq C\varepsilon^{1/2}(N^{-1}\ln N)^{k+1}\quad\text{and}\quad\left\|\eta_{u^{\mathrm{c}}_{1}}\right\|_{\Omega_{22}}\leq C\varepsilon^{1/2}N^{-2\sigma}.$ (3.16)
@@
$\left\|\eta_{u^{\mathrm{b}}_{1}}\right\|_{L^{\infty}(K_{ij})}\leq C\left\|u^{\mathrm{b}}_{1}\right\|_{L^{\infty}(K_{ij})}\leq C\left\|e^{-\beta x/\sqrt{\varepsilon}}\right\|_{L^{\infty}(K_{ij})}\leq Ce^{-\beta\tau/\sqrt{\varepsilon}}\leq CN^{-\sigma}.$ (3.19)
@@
$\left\|\eta_{u^{\mathrm{c}}_{1}}\right\|_{L^{\infty}(K_{ij})}\leq C\left\|u^{\mathrm{c}}_{1}\right\|_{L^{\infty}(K_{ij})}\leq C\left\|e^{-\beta(x+y)/\sqrt{\varepsilon}}\right\|_{L^{\infty}(K_{ij})}\leq Ce^{-\beta\tau/\sqrt{\varepsilon}}\leq CN^{-\sigma}.$ (3.21)
@@ 4 Superapproximation properties of projectors
In Section 4 we present various superapproximation properties for the three types of vertex-edge-element approximations appearing in (3.3).
@@ 4.1 Superapproximation properties of element projector $\Pi$
We prove only the first inequality as the second is derived similarly. For $m=0,1,\dots$, let $\ell_{m}(x)$ be the rescaled Legendre polynomial of degree $m$ defined via Rodrigues’ formula for the standard Legendre polynomial ^$\hat{\ell}_{m}(s)$ on the interval $\hat{I}:=[-1,1]$, namely,
@@
Then one has
^$\left\|\hat{\ell}_{m}\right\|_{\hat{I}}=\sqrt{\frac{2}{2m+1}}\quad\text{ and }\quad\left\|\ell_{m}\right\|_{I_{i}}=\sqrt{\frac{h_{i}}{2m+1}}.$ (4.2)
@@
By (4.4), (4.1), a scaling argument, integration by parts, Cauchy-Schwarz’s inequality and (4.2), one has
@@
$=\frac{1}{\left\|\hat{\ell}_{k+1}\right\|^{2}_{\hat{I}}\left\|\hat{\ell}_{n}\right\|^{2}_{\hat{I}}}\cdot\frac{1}{2^{k+1}(k+1)!}\left|\int_{\hat{K}}\hat{z}(s,t)\partial_{s}^{k+1}(s^{2}-1)^{k+1}\hat{\ell}_{n}(t)\,\mathrm{d}s\,\mathrm{d}t\right|$
$=\frac{2k+3}{2^{k+2}(k+1)!}\cdot\frac{1}{\left\|\hat{\ell}_{n}\right\|^{2}_{\hat{I}}}\left|\int_{\hat{K}}\partial_{s}^{k+1}\hat{z}(s,t)(s^{2}-1)^{k+1}\hat{\ell}_{n}(t)\,\mathrm{d}s\,\mathrm{d}t\right|$ (4.7)
$\leq\frac{1}{\left\|\hat{\ell}_{n}\right\|_{\hat{I}}}\left[\int_{\hat{K}}\left(\partial_{s}^{k+1}\hat{z}\right)^{2}\,\mathrm{d}s\,\mathrm{d}t\right]^{1/2}\leq C\sqrt{2n+1}\left[\frac{1}{|K_{ij}|}\int_{K_{ij}}\left(h_{i}^{k+1}\partial_{x}^{k+1}z\right)^{2}\,\mathrm{d}x\,\mathrm{d}y\right]^{1/2}$
$\leq C\sqrt{2n+1}\sqrt{\frac{h_{i}}{h_{j}}}\,h_{i}^{k}\left\|\partial_{x}^{k+1}z\right\|_{K_{ij}}.$
Here and in what follows $C>0$ is independent of $n,k,h_{i},h_{j}$ and $z$. Similarly,
$|\alpha^{i+1,j}_{k+1,n}|\leq C\sqrt{2n+1}\sqrt{\frac{h_{i+1}}{h_{j}}}\,h_{i+1}^{k}\left\|\partial_{x}^{k+1}z\right\|_{K_{i+1,j}}.$
Our argument now depends on the parity of the nonnegative integer $k$. If $k$ is odd, then $\alpha^{ij}_{k+1,n}+\alpha^{i+1,j}_{k+1,n}(-1)^{k+1}=\alpha^{ij}_{k+1,n}+\alpha^{i+1,j}_{k+1,n}$. Thus, since $h_{i}=h_{i+1}$, the pairwise orthogonality of the rescaled Legendre polynomials, (4.2), Cauchy-Schwarz and inverse inequalities yield
@@
$\leq C\max_{0\leq n\leq k}\Big{(}|\alpha^{ij}_{k+1,n}|+|\alpha^{i+1,j}_{k+1,n}|\Big{)}\left\|\ell_{n}\right\|_{J_{j}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
$\leq C\sqrt{2n+1}\sqrt{\frac{h_{i}}{h_{j}}}\,h_{i}^{k}\left\|\partial_{x}^{k+1}z\right\|_{K_{ij}\cup K_{i+1,j}}\sqrt{\frac{h_{j}}{2n+1}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$